OpenStudy (anonymous):

If ∠PLA and ∠ELA are complementary, ∠PLA = 5x – 2, and ∠ELA = x + 8, what is the measure of ∠PLE?

OpenStudy (anonymous):

give a screenshot of the problem

OpenStudy (anonymous):

Since they're complementary ∠PLA + ∠ELA = 90 So 5x-2+x+8=90 (5x+x-2+8)=6x+6 =90 6x=90-6 6x=84 (By dividing 6 from both sides) x=14 So : ∠PLA = 5(14)-2 = 68

OpenStudy (anonymous):

medal?